OpenStudy (anonymous):

a freight train leaves the train station 4hrs before a passenger train. the two trains are traveling in the same direction on parallel tracks. if the rate of the passenger train is 40 mph faster than the freight train, how fast is each train traveling if the passenger train passes the freight train in 6hrs?

OpenStudy (anonymous):

okay so lets break this down. the Frieght which i will call F leaves 4 hours before the passenger train that i will cal P

OpenStudy (anonymous):

so in six hours how many hours has train P been traveling?

OpenStudy (anonymous):

10?

OpenStudy (anonymous):

okay lets say train F left the station at 6 in the morning. 4 hours later train P left the station it is now 12 noon how many hours has P been traveling?

OpenStudy (aum):

Assume F is the speed of the Freight train and P is the speed of the Passenger train. "passenger train is 40 mph faster than the freight train" P = F + 40 ----- (1) "passenger train passes the freight train in 6hrs" .... "freight train leaves the train station 4hrs before a passenger train" So when the passenger train passes the freight train it has been traveling for 6 hours and the freight train has been traveling for 10 hours and they have traveled the same distance in that time. Therefore, 6 * P = 10 * F ------ (2) Solve for P and F from equations (1) and (2).

OpenStudy (anonymous):

okay thanks